OpenStudy (anonymous):

Can someone just check if im doing this right... Write an equation that expresses the following: w varies jointly with u and d and inversely with p. Use k as the constant of proportionality. I came up with w=k(ud/p) Is that right?

jimthompson5910 (jim_thompson5910):

You are correct saying "w varies jointly with u and d " means w = k*(ud) and saying "w varies...inversely with p" means w = k/p put them together to get what you have \[\large w = k\frac{ud}{p}\]

OpenStudy (anonymous):

yay!! thx @jim_thompson5910 :)

jimthompson5910 (jim_thompson5910):

you're welcome